# Radial Profiles of Radio Halos in Massive Galaxy Clusters: Diffuse Giants Over 2 Mpc

### v1

@@ 1 Introduction
Radio halos are among the largest class of extended diffuse radio sources in the intracluster medium (ICM). They are found at the center of merging clusters and typically correlate with the X-ray emission (e.g., Govoni et al., 2001; Rajpurohit et al., 2018; Botteon et al., 2020; Bonafede et al., 2022). They are believed to be formed by turbulent re-acceleration of cosmic ray electrons (CRe) which become radio-emitting after interacting with turbulence, injected into the ICM by cluster mergers (Brunetti et al., 2001; Petrosian, 2001; Brunetti & Lazarian, 2007; Brunetti & Jones, 2014; Miniati, 2015; van Weeren et al., 2019). Secondary particles, produced via hadronic collisions in the ICM, may provide an additional mechanism to generate CRe in clusters (e.g., Blasi & Colafrancesco, 1999; Dolag & Enßlin, 2000; Pfrommer et al., 2008). However, the current gamma-ray limits disfavor this scenario (Pinzke et al., 2017; Brunetti et al., 2017).
@@
The remarkable developments in radio sensitivity and resolution of modern instruments are uncovering larger extents of these diffuse radio halos as well as intricate structures in their emission (e.g., Knowles et al., 2022; Botteon et al., 2022a, b; Duchesne et al., 2024; Botteon et al., 2024; Rajpurohit et al., 2023, 2021a; van Weeren et al., 2017, 2016; Osinga et al., 2024; Sikhosana et al., 2023). In particular, while a few radio halos with very large linear size (LLS), exceeding 2 Mpc, were already reported with “old generation” interferometers (e.g. Bullet and Abell 2163 Feretti et al., 2001; Liang et al., 2000). Radio halos with such extents are nowadays detected in increasing number of systems, see Figure 1 for the complete list (Sikhosana et al., 2023; Bruno et al., 2023; Rajpurohit et al., 2021b, a; Bonafede et al., 2022; Shweta et al., 2020). With the deepest low-frequency observations carried out on a cluster with LOFAR, for example, a total extent of approximately 5 Mpc for the diffuse emission in Abell 2255 has been found (Botteon et al., 2022b). Additionally, even at high frequencies, diffuse emission, extending approximately 6 Mpc, is observed in PLCK  G287.0+32.9 (Rajpurohit et al. in prep.).
@@
Recently, by analyzing a sample of $\sim 310$ Planck clusters, Cuciti et al. (2022) suggested the existence of a new class of radio sources, called mega-halos, based on the detection of four cluster halos that exhibit properties distinct from those typically observed in classical radio halos. These sources were detected with LOFAR (at low resolution) and were found to be hosted in massive galaxy clusters (see Figure 1). Their defining characteristics include a LLS exceeding 2 Mpc, extending at least up to $R_{500}$ (the radius within which the average mass density of a cluster is 500 times the critical density of the Universe), and a two-component radial radio surface brightness profile. The first component appears as a brighter, canonical halo-like structure following an exponential profile, while the second component is reported to be shallower. Additionally, based on the spectral analysis performed at 50-150 MHz for two clusters, the outer component of the mega-halo is reported to exhibit an ultra-steep spectrum ($\alpha\geq-1.5$)
@@
The presence of two components in the radial surface brightness profiles has also been observed in relaxed clusters hosting mini-halo (related to sloshing motions) and halo like emission (e.g., Biava et al., 2024; van Weeren et al., 2024), dubbed hybrid halos, which typically shows LLS below 1 Mpc. In these systems, the outer, fainter component exhibits characteristics typical of classical halos. Moreover, some classical radio halos also exhibit multiple components within the halo emission (Rajpurohit et al., 2021a; Bruno et al., 2023; Rajpurohit et al., 2023). Whether these two-component systems (halos and mega-halos) constitute a new class of sources that lie below or encompass classical radio halos remains an open question, requiring deeper multi-frequency observations and larger sample studies.
@@
The main goal of this paper is to investigate the observational properties of radio halos, where the new observations reveal additional extents of their diffuse emission. Using high quality observations, we analyzed radial profiles of the following five clusters: PLCK G287.0+32.9, Abell 2744, the Bullet Cluster, MACS J0717+3745, and Abell 2142. These targets were selected due to their high mass and availability of deep multifrequency radio observations. We note that all reported mega-halos are detected in the redshift range of $z\sim 0.17-0.28$ and mass $M_{500}\sim 6-11\times 10^{14}M_{\odot}$. Therefore, PLCK G287+32.9, Abell 2744, the Bullet Cluster, MACS J0717+3745, and Abell 2142 are ideal clusters to host mega-halos as supported by their location in the $M-z$ distribution (see Figure 1). We investigate whether or not these halos with large extents exhibit characteristics typical of classical halos.
@@
Figure 1: Mass-redshift distribution of the PSZ2 clusters. The subsample of PSZ2 in LoTSS-DR2 (Botteon et al., 2022a) is reported in black. Circles denote clusters with radio halos having LLS $>$ 2 Mpc and are color-coded as follows: in cyan are the mega-halos (Cuciti et al., 2022), in red are the clusters analyzed in this work, and in magenta are other clusters reported in the literature. The solid lines represent the boundaries above which mega halos are expected to be detected with current LOFAR observations assuming that their surface brightness is $\propto M_{500}^{\beta}(1+z)^{-4}$, for three possibilities for $\beta$.
@@
Throughout this paper, we adopt a flat $\Lambda$CDM cosmology with $H_{\rm{0}}=70$ km s-1 Mpc-1, $\Omega_{\rm{m}}=0.30$, and $\Omega_{\Lambda}=0.71$. We define the spectral index, $\alpha$, so that $S_{\nu}\propto\nu^{\alpha}$, where $S$ is the flux density at frequency $\nu$ ($\alpha\leq-1$). All output radio images are in the J2000 coordinate system. The LLS and radial profiles of all radio halos presented in this study are measured at $\geq 3\sigma_{\rm rms}$ level.
@@ 2 Observations and data reduction
The cluster was observed with MeerKAT in the UHF 0.55-1.0 GHz (Project code: SCI-20220822), L-band 1-1.7 GHz (Project code: SCI-20220822) and S-band 1.9-2.8 GHz (Project code: SCI-20230907). All four polarization products were recorded using the 4K correlator mode, covering a total bandwidth 544 MHz and 875 MHz at UHF and S-band, respectively. For the UHF and S-band observations, J0408-6545 was the primary calibrator used for flux and bandpass calibration, observed at the beginning and end of the observing run. J1154-3505 was observed as a gain calibrator and J0521+1638 as a polarization calibrator.
@@
After initial calibration, we created an initial image of the target field using WSClean (Offringa et al., 2014) within CARACAL. Three rounds of phase-only self-calibration were performed using CubiCal, followed by a final round of amplitude-phase calibration. The calibrated data were imaged using the Briggs weighting scheme with a robust parameter of 00, and multiscale cleaning. For the L-band data reduction, we refer to Balboni et al. in prep.
We also observed the cluster with the upgraded Giant Metrewave Radio Telescope (uGMRT) at Band3 (Project code: 45_020 ), covering the frequency range of 300-500 MHz. The uGMRT data calibration was performed using the Source Peeling and Atmospheric Modeling pipeline (SPAM; Intema et al., 2009). We first split the wideband dataset into six sub-bands. The flux density of the primary calibrator 3C 286 was set according to Scaife & Heald (2012). Following this, the data were averaged, flagged, and corrected for the bandpass. We used a global sky model obtained from the GMRT GSB data to correct the phase gains of the target. Finally, the SPAM calibrated sub-bands were imaged in WSClean to produce deep, full continuum images using Briggs weighting scheme with a robust parameter of 00, and multiscale cleaning.
Figure 2: PLCK G287.0+32.9 field (square-root scale) full band continuum images from MeerKAT observations using S-band (1.9-2.8 GHz), L-band (0.9-1.7 GHz) and UHF bands (0.5-1.0 GHz) and uGMRT Band3 (300-500 MHz). All images have a common resolution of 20″. The radio beam size is indicated in the bottom left corner of each image. The images show large scale diffuse radio emission including two symmetrically located relics and a 3.5 Mpc radio halo. The radio maps were created using a Briggs weighing scheme with robust parameter of $-0.5$ and without any uv-cut. The rms noise levels are $\rm\sigma_{2.4\,GHz}=8\mu\,Jy\,beam^{-1}$, $\rm\sigma_{1.28\,GHz}=15\mu\,Jy\,beam^{-1}$, $\rm\sigma_{815\,MHz}=25\mu\,Jy\,beam^{-1}$, and $\rm\sigma_{350\,MHz}=32\mu\,Jy\,beam^{-1}$
Figure 3: A2744 full band continuum images of the central field from MeerKAT UHF (0.55-1 GHz) and uGMRT Band4 (550-850 MHz) observations (in square-root scale). The figures show central halo emission and four relics in the outskirts (Rajpurohit et al., 2021a). Both images have a common resolution of 25″. The radio beam size is indicated in the bottom left corner of each image. The radio maps were created using a Briggs weighing scheme with robust parameter of 00 and without any uv-cut. The rms noise levels are $\rm\sigma_{815\,MHz}=20\mu\,Jy\,beam^{-1}$ and $\rm\sigma_{675\,MHz}=25\mu\,Jy\,beam^{-1}$.
@@ 2.2 Abell 2744
For the uGMRT Band4 data reduction, we refer to Rajpurohit et al. (2021a), as we used the same calibrated data presented in that work. Both MeerKAT and uGMRT calibrated data were imaged in WSClean using Briggs weighting scheme with a robust parameter of 00, and multiscale cleaning.
@@ 2.3 Bullet cluster
We used calibrated MeerKAT L-band data presented in Sikhosana et al. (2023). For the UHF-band, we used new MeerKAT observations (Project code: SCI-20241101). The observations were performed over two observing runs. J0408-6545 and J1939-6342 were used as flux calibrators, while J0825-5010 as gain calibrator. The observations detail are summarized in Table 3. The data were processed by the SARAO Science Data Processor pipeline [^2] https://skaafrica.atlassian.net/wiki/spaces/ESDKB/pages/338723406/SDP+pipelines+overview. The calibrated data were imaged in WSClean using Briggs weighting scheme with a robust parameter of 00, and multiscale cleaning.
@@
Figure 4: Bullet cluster MeerKAT L-band (0.9-1.7 GHz) and UHF (0.55-1.0 GHz) full band continuum radio images in square-root scale, showing a large scale central halo emission and a toothbrush shaped relic to the east (Sikhosana et al., 2023). Both images have a common resolution of 15″. The radio beam size is indicated in the bottom left corner of each image. The images show large scale asymmetrical halo emission. The radio maps were created using a Briggs weighing scheme with robust parameter of $-0.5$ and without any uv-cut. The rms noise levels are $\rm\sigma_{1.28\,GHz}=18\mu\,Jy\,beam^{-1}$ and $\rm\sigma_{815\,MHz}=26\mu\,Jy\,beam^{-1}$.
@@ 3 Radio continuum maps of individual clusters
Figure 2 shows our new MeerKAT (UHF, L and S bands) and uGMRT Band3 images of PLCK G287+32.9. At a redshift of $z=0.39$, this cluster stands out as an exceptionally luminous and massive (the second most massive) galaxy cluster in the Planck sample (Planck Collaboration et al., 2016). A pair of radio relics, a central radio halo, and three filamentary features were previously reported in the GMRT, Very Large Array, and Murchison Widefield Array observations (Bagchi et al., 2011; Bonafede et al., 2014; George et al., 2017). Our new radio observations revealed 6 Mpc diffuse radio emission filling the entire cluster volume. In this paper, we only focus on the radio halo emission.
We find that the halo in PLCK G287+32.9 is about 3, 3.5, 3.5, and 2.5 Mpc in diameter at 350 MHz, 815 MHz, 1.28 GHz, and 2.4 GHz, respectively (Figure 2). To the best of our knowledge, this is the first halo with LLS is about 2.5 Mpc at 2.4 GHz. The total extent of the PLCK G287+32.9 halo is at least 2.5 times larger than reported earlier (Bonafede et al., 2014). The overall morphology of the halo emission is similar from 350 MHz to 2.4 GHz. However, the halo is more extended toward low frequencies as found in other well-studied halos, for example 1RXS J0603.3+4214 (van Weeren et al., 2016; Rajpurohit et al., 2020), Abell 2744 (Pearce et al., 2017; Rajpurohit et al., 2021a), MACSJ0717+3745 (van Weeren et al., 2017; Bonafede et al., 2018; Rajpurohit et al., 2021b), CIZA J2242.8+5301 (Di Gennaro et al., 2018; Hoang et al., 2017). At 350 MHz, we do not recover the faint outermost regions of the PLCK G287+32.9 halo emission observed in the MeerKAT UHF/L-band. This is due to the superior sensitivity and dense inner uv coverage of the MeerKAT compared to uGMRT.
Figure 5: MAC J0717 LOFAR HBA (120-167 MHz) and Abell 2142 (120-167 MHz) full band continuum radio images in square-root scale. The central region of MAC J0717 is dominated by a bright relic, projected on the the halo emission (Rajpurohit et al., 2021b). The halo in Abell 2142 is extended along the northwest to southeast direction (Bruno et al., 2023). To the north is a tailed radio galaxy (extending E-W). The radio beam size is indicated in the bottom left corner of each image. The rms noise levels are $\rm\sigma_{144\,MHz}=240\mu\,Jy\,beam^{-1}$ for MACS J0717+3745 and $\rm\sigma_{144\,MHz}=400\mu\,Jy\,beam^{-1}$ for Abell 2142.
@@
The new MeerKAT UHF and previously published uGMRT Band 4 images of Abell 2744 are shown in Figure 3. At $z=0.306$, this highly disturbed cluster is very rich at radio, X-ray, and optical wavelengths (Govoni et al., 2001; Owers et al., 2011; Merten et al., 2011; Kempner & David, 2004; Golovich et al., 2019; Jauzac et al., 2018; Venturi et al., 2013; Orrú et al., 2007; George et al., 2017; Paul et al., 2019). It hosts a giant radio halo at its center, a 1.5 Mpc long relic to the northeast, and three additional fainter relics (Pearce et al., 2017; Rajpurohit et al., 2021a). The LLS of the Abell 2744 halo is 2.6 Mpc at 675 MHz, exhibiting a roughly circular morphology. In our MeerKAT UHF map (0.55-1.0 GHz), we recovered the full 2.6 Mpc extent of the radio halo emission along with other known features (Figure 3). At 1.5 GHz, the halo extends to 2.1 Mpc, suggesting a steep spectrum in the outer regions (Pearce et al., 2017). By comparing the point-to-point radio brightness and spectral index with the X-ray brightness and temperature distribution, Rajpurohit et al. (2021a) found that the Abell 2744 halo consists of multiple components.
@@
Figure 6: Concentric annuli and sectors used to extract the radial brightness profiles of the radio halos in PLCK G287.0+32.9 (50″  resolution), Abell 2744 (25″  resolution), Bullet (15″  resolution), MACS J0717.5+3745 (20″  resolution), and Abell 2142 (75″  resolution; image adopted from Bruno et al., 2023). White pataches are the masked regions. The width of each circular annulus is half the full width at half maximum of the beam size. The last panel illustrates the different azimuthal sectors (red elliptical ones) used to investigate the radial profiles of the PLCK G287+32.9 halo, with the resulting profiles shown in Figure 11.
@@
Figure 4 shows the MeerKAT UHF and L-band images of the Bullet Cluster. The cluster is known to host a radio halo with LLS  2.4 Mpc and a toothbrush shaped relic to the east (Sikhosana et al., 2023; Shimwell et al., 2014). In our new UHF image, the full extent of the halo emission, as seen at L-band, is recovered. The halo is asymmetric, extending primarily in the north-south and east-west directions.
@@
Figure 5 shows the low-resolution LOw- Frequency ARray (LOFAR) High-Band Antenna (HBA) images of the halos in MACS J0717+3745 (left panel: Rajpurohit et al., 2021b), and Abell 2142 (right panel: Bruno et al., 2023), with LLS of 2.6 Mpc and 2.5 Mpc, respectively. Similar to PLCK G287+32.9, both halos are asymmetrical, particularly MACS J0717+3745, which is known to host a bright filamentary relic at its center along with other linear filaments (van Weeren et al., 2017; Rajpurohit et al., 2021c). The LLS of the halos in MACS J0717+3745 and Abell 2142 is less than 2 Mpc at frequencies > 650 MHz (van Weeren et al., 2017; Rajpurohit et al., 2021b; Riseley et al., 2024).

### v2

@@ 1 Introduction
Radio halos are among the largest class of extended diffuse radio sources in the intracluster medium (ICM). They are found at the center of merging clusters and typically correlate with the X-ray emission (e.g., Govoni et al., 2001; Rajpurohit et al., 2018; Botteon et al., 2020; Bonafede et al., 2022). They are believed to be formed by turbulent re-acceleration of cosmic ray electrons (CRe) which become radio-emitting by turbulence, injected into the ICM by cluster mergers (Brunetti et al., 2001; Petrosian, 2001; Brunetti & Lazarian, 2007; Brunetti & Jones, 2014; Miniati, 2015; van Weeren et al., 2019). Secondary particles, produced via hadronic collisions in the ICM, may provide an additional mechanism to generate CRe in clusters (e.g., Blasi & Colafrancesco, 1999; Dolag & Enßlin, 2000; Pfrommer et al., 2008). However, the current gamma-ray limits disfavor this scenario (Pinzke et al., 2017; Brunetti et al., 2017).
@@
The remarkable developments in radio sensitivity and resolution of modern instruments are uncovering larger extents of these diffuse radio halos along with intricate embedded structures (e.g., Knowles et al., 2022; Botteon et al., 2022a, b; Duchesne et al., 2024; Botteon et al., 2024; Rajpurohit et al., 2023, 2021a; van Weeren et al., 2017, 2016; Osinga et al., 2024; Sikhosana et al., 2023). In particular, while a few radio halos with very large linear size (LLS), exceeding 2 Mpc, were already reported with “old generation” interferometers (e.g. Bullet and Abell 2163 Feretti et al., 2001; Liang et al., 2000). Radio halos with such extents are being detected more frequently in a growing number of systems, see Figure 1 (Sikhosana et al., 2023; Bruno et al., 2023; Rajpurohit et al., 2021b, a; Bonafede et al., 2022; Shweta et al., 2020). With the deepest low-frequency observations carried out on a cluster with LOFAR, for example, a total extent of approximately 5 Mpc for the diffuse emission in Abell 2255 has been found (Botteon et al., 2022b). Additionally, even at high frequencies, diffuse emission, extending approximately 6 Mpc, is observed in PLCK  G287.0+32.9 (Rajpurohit et al. in prep.).
@@
Recently, by analyzing a sample of $\sim 310$ Planck clusters, Cuciti et al. (2022) suggested the existence of a new class of radio sources, called mega-halos, based on the detection of diffuse emission in four clusters that exhibit properties distinct from those typically observed in classical radio halos. These sources were detected with LOFAR (at low resolution) and were found to be hosted in massive galaxy clusters (see Figure 1). Their defining characteristics include an LLS exceeding 2 Mpc, extending at least up to $R_{500}$ (the radius within which the average mass density of a cluster is 500 times the critical density of the Universe), and a two-component radial radio surface brightness profile. The first component appears as a brighter, canonical halo-like structure following an exponential profile, while the second component is reported to be shallower. Additionally, based on the spectral analysis performed at 50-150 MHz for two clusters, the outer component of the mega-halo is reported to exhibit an ultra-steep spectrum ($\alpha\geq-1.5$)
@@
The presence of two components in the radial surface brightness profiles has also been observed in relaxed clusters hosting mini-halo (related to sloshing motions) and halo like emission (e.g., Biava et al., 2024; van Weeren et al., 2024), dubbed hybrid halos, which typically show LLS below 1 Mpc. In these systems, the outer, fainter component exhibits characteristics typical of classical halos. Moreover, some classical radio halos also exhibit multiple components within the halo emission (Rajpurohit et al., 2021a; Bruno et al., 2023; Rajpurohit et al., 2023). Whether these two-component systems (halos and mega-halos) constitute a new class of sources that lie below or encompass classical radio halos remains an open question, requiring deeper multi-frequency observations and larger sample studies.
@@
The main goal of this paper is to investigate the observational properties of radio halos, where the new observations reveal additional extents of their diffuse emission. Using high-quality observations, we analyzed radial profiles of the following five clusters: PLCK G287.0+32.9, Abell 2744, the Bullet Cluster, MACS J0717+3745, and Abell 2142. These targets were selected due to their high mass and availability of deep, multifrequency radio observations. We note that to date mega-halos are detected in the redshift range of $z\sim 0.17-0.28$ and mass $M_{500}\sim 6-11\times 10^{14}M_{\odot}$. Their surface brightness is assumed to scale with $\propto M_{500}^{\beta}(1+z)^{-4}$. In Figure 1, we show the three possibilities for $\beta$. PLCK G287+32.9, Abell 2744, the Bullet Cluster, MACS J0717+3745, and Abell 2142 are ideal clusters to host mega-halos as supported by their location in the MzM-z distribution. We investigate whether or not these halos with large extents exhibit characteristics typical of classical halos.
@@
Figure 1: Cluster mass versus redshift distribution of the PSZ2 clusters. The subsample of PSZ2 in LoTSS-DR2 (Botteon et al., 2022a) is reported in black. Circles denote clusters with halos having LLS >> 2 Mpc and are color-coded as follows: in cyan are the mega-halos (Cuciti et al., 2022), in red are the clusters analyzed in this work, and in magenta are other clusters reported in the literature. The solid lines represent the boundaries above which mega-halos are assumed to be detected with LOFAR observations assuming that their surface brightness is $\propto M_{500}^{\beta}(1+z)^{-4}$, for three possibilities for $\beta$.
@@
[…]
Table 4: Imaging properties of radio maps used in the analysis
Cluster Central Frequency Image Name Restoring Beam Robust uv-cut uv-taper RMS noise
GHz parameter $\upmu\rm Jy\,beam^{-1}$
PLCK G287.0+32.9 2.40 IM1 $20\arcsec\times 20\arcsec$ $0.0$ - $15\arcsec$ 8
1.28 IM2 $20\arcsec\times 20\arcsec$ $0.0$ - 15″ 15
0.815 IM3 $20\arcsec\times 20\arcsec$ $0.0$ - 15″ 25
0.400 IM4 $20\arcsec\times 20\arcsec$ $0.0$ - 15″ 32
[…]
Abell 2744 0.815 IM9 $25\arcsec\times 20\arcsec$ $0.0$ - $20\arcsec$ 24
0.675 IM10 $25\arcsec\times 25\arcsec$ $0.0$ - 20″ 20
[…]
Bullet 1.28 IM14 $20\arcsec\times 20\arcsec$ $0.0$ - $15\arcsec$ 18
0.815 IM15 $20\arcsec\times 20\arcsec$ $0.0$ - 15″ 20
[…]
Note. — Final imaging was performed in WSCLEAN using multiscale and with the Briggs weighting scheme. For the Abell 2142 image properties, we refer to Bruno et al. (2023)
@@ 2 Observations and data reduction
The cluster was observed with MeerKAT in the UHF 0.55-1.0 GHz (Project code: SCI-20220822), L-band 1-1.7 GHz (Project code: SCI-20220822) and S-band 1.9-2.8 GHz (Project code: SCI-20230907), see Table 1 for details. All four polarization products were recorded using the 4K correlator mode, covering a total bandwidth 544 MHz and 875 MHz at UHF and S-band, respectively. For the UHF and S-band observations, J0408-6545 was the primary calibrator used for flux and bandpass calibration, observed at the beginning and end of the observing run. J1154-3505 was observed as a gain calibrator and J0521+1638 as a polarization calibrator.
@@
After initial calibration, we flagged the remaining low surface brightness RFI using AOflagger. We then created an initial image of the target field using WSClean (Offringa et al., 2014) within CARACAL. Three rounds of phase-only self-calibration were performed using CubiCal (Kenyon et al., 2018), followed by a final round of amplitude-phase calibration. The calibrated data were imaged in WSClean using the Briggs weighting scheme with a robust parameter of 0, and multiscale cleaning. For the L-band data reduction, we refer to Balboni et al. in prep.
We also observed the cluster with the upgraded Giant Metrewave Radio Telescope (uGMRT) at Band3 (Project code: 45_020 ), covering the frequency range of 300-500 MHz. The uGMRT data calibration was performed using the Source Peeling and Atmospheric Modeling pipeline (SPAM; Intema et al., 2009). We first split the wideband dataset into six sub-bands. The flux density of the primary calibrator 3C 286 was set according to Scaife & Heald (2012). Following this, the data were averaged, flagged, and corrected for the bandpass. We used a global sky model obtained from the GMRT GSB data to correct the phase gains of the target. Finally, the SPAM calibrated sub-bands were imaged in WSClean to produce deep, full continuum images using Briggs weighting scheme with a robust parameter of 0, and multiscale cleaning.
Figure 2: PLCK G287.0+32.9 field (square-root scale) full band continuum images from MeerKAT observations using S-band (1.9-2.8 GHz), L-band (0.9-1.7 GHz) and UHF bands (0.5-1.0 GHz) and uGMRT Band3 (300-500 MHz). All images have a common resolution of 20″. The radio beam size is indicated in the bottom left corner of each image. The images show large scale diffuse radio emission including two symmetrically located relics and a 3.5 Mpc radio halo. The contour levels are drawn at $1,2,4,8...]\times 3\sigma_{\rm rms}$. Dashed contours depict the $3\sigma_{\rm rms}$ contours. For image properties, see Table 4 (IM1, IM2, IM3 and IM4).
Figure 3: PLCK G287.0+32.9 field (square-root scale) full band continuum images from MeerKAT observations using S-band (1.9-2.8 GHz), L-band (0.9-1.7 GHz) and UHF bands (0.5-1.0 GHz) and uGMRT Band3 (300-500 MHz). All images have a common resolution of 50″. The radio beam size is indicated in the bottom left corner of each image. The images are used to extract radial surface brightness profiles of the halo as a function of frequency. The contour levels are drawn at $1,2,4,8...]\times 3\sigma_{\rm rms}$. Dashed contours depict the $3\sigma_{\rm rms}$ contours. For image properties, see Table 4 (IM5, IM6, IM7 and IM8).
Figure 4: A2744 full band continuum images of the central field from MeerKAT UHF (0.55-1 GHz) and uGMRT Band4 (550-850 MHz) observations (in square-root scale). The figures show central halo emission and four relics in the outskirts (Rajpurohit et al., 2021a). Both images have a common resolution of 25″. The radio beam size is indicated in the bottom left corner of each image. The contour levels are drawn at $1,2,4,8...]\times 3\sigma_{\rm rms}$. Dashed contours depict the $3\sigma_{\rm rms}$ contours. For image properties, see Table 4 (IM9 and IM10).
@@ 2.2 Abell 2744
For the uGMRT Band4 data reduction, we refer to Rajpurohit et al. (2021a), as we used the same calibrated data presented in that work. Both MeerKAT and uGMRT calibrated data were imaged in WSClean using Briggs weighting scheme with a robust parameter of 0, and multiscale cleaning.
@@ 2.3 Bullet cluster
We used calibrated MeerKAT L-band data presented in Sikhosana et al. (2023). For the UHF-band, we used new MeerKAT observations (Project code: SCI-20241101). The observations were performed over two observing runs. J0408-6545 and J1939-6342 were used as flux calibrators, while J0825-5010 as a gain calibrator. The observational details are summarized in Table 3. The data were processed by the SARAO Science Data Processor pipeline [^2] https://skaafrica.atlassian.net/wiki/spaces/ESDKB/pages/338723406/SDP+pipelines+overview. The calibrated data were imaged in WSClean using Briggs weighting scheme with a robust parameter of 0, and multiscale cleaning.
@@
Figure 5: Bullet cluster MeerKAT L-band (0.9-1.7 GHz) and UHF (0.55-1.0 GHz) full band continuum radio images in square-root scale, showing a large scale central halo emission and a toothbrush shaped relic to the east (Sikhosana et al., 2023). Both images have a common resolution of 15″. The radio beam size is indicated in the bottom left corner of each image. The images show large scale asymmetrical halo emission. The contour levels are drawn at $1,2,4,8...]\times 3\sigma_{\rm rms}$. Dashed contours depict the $3\sigma_{\rm rms}$ contours. For image properties, see Table 4 (IM14 and IM15).
@@ 3 Radio continuum maps of individual clusters
Figures 2 and 3 show our new MeerKAT (UHF, L and S bands) and uGMRT Band3 images of PLCK G287+32.9. At a redshift of $z=0.39$, this cluster stands out as an exceptionally luminous and massive (the second most massive) in the Planck sample (Planck Collaboration et al., 2016). A pair of radio relics, a central radio halo, and three filamentary features were previously reported in the GMRT, Very Large Array, and Murchison Widefield Array observations (Bagchi et al., 2011; Bonafede et al., 2014; George et al., 2017). Our new radio observations revealed 6 Mpc diffuse radio emission filling the entire cluster volume and a multitude of new structures. In this paper, we only focus on the radio halo emission.
We find that the halo in PLCK G287+32.9 is about 3, 3.5, 3.5, and 2.5 Mpc in diameter at 350 MHz, 815 MHz, 1.28 GHz, and 2.4 GHz, respectively (Figure 3). To the best of our knowledge, this is the first halo with LLS of about 2.5 Mpc at 2.4 GHz. The total extent of the PLCK G287+32.9 halo is at least 2.5 times larger than reported earlier (Bonafede et al., 2014). The overall morphology of the halo emission is similar from 350 MHz to 2.4 GHz. However, the halo is more extended toward low frequencies as found in other well-studied halos, for example, 1RXS J0603.3+4214 (van Weeren et al., 2016; Rajpurohit et al., 2020), Abell 2744 (Pearce et al., 2017; Rajpurohit et al., 2021a), MACSJ0717+3745 (van Weeren et al., 2017; Bonafede et al., 2018; Rajpurohit et al., 2021b), CIZA J2242.8+5301 (Di Gennaro et al., 2018; Hoang et al., 2017). At 350 MHz, we do not recover the faint outermost regions of the PLCK G287+32.9 halo observed in the MeerKAT UHF/L-band images. This is due to the superior sensitivity and dense inner uv-coverage of the MeerKAT compared to the uGMRT.
Figure 6: MAC J0717 LOFAR HBA (120-167 MHz) and Abell 2142 (120-167 MHz) full band continuum radio images in square-root scale. The central region of MAC J0717 is dominated by a bright relic, projected on the halo emission (Rajpurohit et al., 2021b). The halo in Abell 2142 is extended along the northwest to southeast direction (Bruno et al., 2023). To the north is a tailed radio galaxy (extending E-W). The radio beam size is indicated in the bottom left corner of each image. The contour levels are drawn at $1,2,4,8...]\times 3\sigma_{\rm rms}$. Dashed contours depict the $3\sigma_{\rm rms}$ contours. For MACS J0717+3745 image properties, see Table 4 (IM18 and IM19). The noise level of the LOFAR image of Abell 2142 image is  $\rm\sigma_{144\,MHz}=400\mu\,Jy\,beam^{-1}$.
@@
The new MeerKAT UHF and previously published uGMRT Band 4 images of Abell 2744 are shown in Figure 4. Located at a redshift of $z=0.306$, this highly disturbed cluster is very rich in radio, X-ray, and optical wavelengths (Govoni et al., 2001; Owers et al., 2011; Merten et al., 2011; Kempner & David, 2004; Golovich et al., 2019; Jauzac et al., 2018; Venturi et al., 2013; Orrú et al., 2007; George et al., 2017; Paul et al., 2019). It is known to host a giant radio halo at its center, a 1.5 Mpc long relic to the northeast, and three additional fainter relics (Pearce et al., 2017; Rajpurohit et al., 2021a; Knowles et al., 2022). The LLS of the Abell 2744 halo is 2.6 Mpc at 675 MHz, exhibiting a roughly circular morphology. In our MeerKAT UHF map (0.55-1.0 GHz), we recovered the full 2.6 Mpc extent of the radio halo emission along with other known features (Figure 4). At 1.5 GHz, the halo extends to 2.1 Mpc, suggesting a steep spectrum in the outer regions (Pearce et al., 2017). By comparing the point-to-point radio brightness and spectral index with the X-ray brightness and temperature distribution, Rajpurohit et al. (2021a) found that the Abell 2744 halo consists of multiple components.
@@
Figure 7: Concentric annuli used to extract the radial brightness profiles of the radio halos in PLCK G287.0+32.9 (50″  resolution), Abell 2744 (25″  resolution), Bullet (15″  resolution), MACS J0717.5+3745 (20″  resolution), and Abell 2142 (75″  resolution; image adopted from Bruno et al., 2023). The hatched areas indicate the masked regions. The width of each circular annulus is half the full width at half maximum of the beam size. The last panel illustrates the different azimuthal annuli (elliptical and partial elliptical) used to investigate the radial profiles of the PLCK G287+32.9 halo, with the resulting profiles shown in Figure 12 top right panel. The last panel image of PLCK G287.0+32.9 is obtained by subtracting discrete sources using an inner uvcut. For the rest of the image properties, see Table 4 (IM7, IM12, IM15, and IM19). We emphasize that the masked regions differ between the PLCK G287.0+32.9 maps with masking and those with uv-subtraction (see first and last panel images).
@@
Figure 5 shows the MeerKAT UHF and L-band images of the Bullet Cluster. The cluster is known to host a radio halo with LLS of about 2.4 Mpc and a toothbrush-shaped relic to the east (Sikhosana et al., 2023; Shimwell et al., 2014). In our new UHF image, the full extent of the halo emission reported at L-band is recovered. The halo is asymmetric, extending primarily in the north-south and east-west directions.
@@
Figure 6 shows the low-resolution LOw- Frequency ARray (LOFAR) High-Band Antenna (HBA) images of the halos in MACS J0717+3745 (left panel: Rajpurohit et al., 2021b), and Abell 2142 (right panel: Bruno et al., 2023), with LLS of about 2.6 Mpc and 2.5 Mpc, respectively. Similar to PLCK G287+32.9, both of these halos are asymmetrical, particularly MACS J0717+3745, which is known to host a bright filamentary relic at its center along with other linear filaments (van Weeren et al., 2017; Rajpurohit et al., 2021c). The LLS of the halos in MACS J0717+3745 and Abell 2142 is less than 2 Mpc at frequencies >650>650 MHz (van Weeren et al., 2017; Rajpurohit et al., 2021b; Riseley et al., 2024).